# NEST: Self-supervised Fast Conformer as All-purpose Seasoning to Speech Processing Tasks

### v4

@@ Abstract
Self-supervised learning (SSL) has been proved to benefit a wide range of speech processing tasks, such as speech recognition/translation, speaker verification and diarization, etc. However, most of current speech SSL approaches are computationally expensive. In this paper, we introduce a simplified and more efficient SSL framework, termed as NeMo Encoder for Speech Tasks (NEST). Specifically, we adopt the FastConformer architecture with 8x sub-sampling rate, which is faster than Transformer or Conformer architectures. Instead of clustering-based quantization, we use fixed random projection for its simplicity and effectiveness. We also implement a generalized noisy speech augmentation that teaches the model to disentangle the main speaker from noise or other speakers. Experiments show that NEST improves over existing self-supervised models and achieves new state-of-the-art performance on a variety of speech processing tasks, such as speech recognition/translation, speaker diarization, spoken language understanding, etc. Code and checkpoints will be publicly available via NVIDIA NeMo framework [^1] https://github.com/NVIDIA/NeMo.
@@ II-B Speech Augmentation
TABLE II: Results on multi-lingual ASR with punctuation and capitalization. Performance is evaluated by word error rate (WER) including native punctuation and capitalization from the source datasets.
@@ III-B Results on SUPERB Multi-task Speech Processing
[1.5ex] NEST-XL-Transformer 1B 42K 30.87 39.95 30.01 22.82 34.92 27.99 29.50 22.61 39.27 27.73 32.51 32.42
TABLE IV: DER results on speaker diarization. Underline indicates the second best evalutations. Starred(*) systems are not end-to-end systems which involve clustering steps.
@@
RandFC-L-MLP 21.71 11.60 15.89 24.34
NEST-L-MLP 16.83 7.88 11.71 23.46
RandFC-L-Sortformer-HL [37] 18.93 9.39 13.56 23.30
NEST-L-Sortformer-HL [37] 16.06 6.40 11.02 20.10
@@ III-E Results on Speaker Diarization
To assess the impact of NEST on speaker diarization, we train two variants of end-to-end diarization models: (1) simple two-layer multi-layer-perceptron (MLP) on top of FastConformer encoder and train with PIL; (2) more sophisticated Sortformer [37] hybrid loss (HL) model with post processing (PP) [^2] Post-processing parameters were tuned separately for DIHARD3 and CALLHOME on corresponding training parts. with 18 layers of transformer on top of the encoder. We also apply NEST and random initialization to both models for comparison. For training data, we use a combination of 2030 hours of real data (Fisher English [24], AMI Mix-Headset train+dev [48], ICSI [49], DIHARD3 dev [50], VoxConverse v0.3 [51], AISHELL-4 [52], CALLHOME-part1333We follow splits from the Kaldi x-vector recipe [53] by using part1 for training and part2 for evaluation. [54]) and 5150 hours of simulated data (composed from LibriSpeech [43] and SRE [55, 56]) generated by the NeMo speech data simulator [57]. We evaluate models’ performance on DIHARD3-eval [50] and CALLHOME-part2 [54].

### v5

@@ Abstract
Self-supervised learning (SSL) has been proved to benefit a wide range of speech processing tasks, such as speech recognition/translation, speaker verification and diarization, etc. However, most of current speech SSL approaches are computationally expensive. In this paper, we introduce a simplified and more efficient SSL framework, termed as NeMo Encoder for Speech Tasks (NEST). Specifically, we adopt the FastConformer architecture with 8x sub-sampling rate, which is faster than Transformer or Conformer architectures. Instead of clustering-based quantization, we use fixed random projection for its simplicity and effectiveness. We also implement a generalized noisy speech augmentation that teaches the model to disentangle the main speaker from noise or other speakers. Experiments show that NEST improves over existing self-supervised models and achieves new state-of-the-art performance on a variety of speech processing tasks, such as speech recognition/translation, speaker diarization, spoken language understanding, etc. Code and checkpoints are publicly available via NVIDIA NeMo framework [^1] https://github.com/NVIDIA/NeMo [^2] https://huggingface.co/nvidia/ssl_en_nest_large_v1.0333https://huggingface.co/nvidia/ssl_en_nest_xlarge_v1.0.
@@ II-B Speech Augmentation
TABLE II: Results on multi-lingual ASR with punctuation and capitalization. Performance is evaluated by word error rate (WER) including native punctuation and capitalization from the source datasets. Underline indicates the second best performance.
@@ III-B Results on SUPERB Multi-task Speech Processing
NEST-XL-Transformer 1B 42K 30.87 39.95 30.01 22.82 34.92 27.99 29.50 22.61 39.27 27.73 32.51 32.42
TABLE IV: DER results on speaker diarization. Underline indicates the second best performance. Starred(*) systems are not end-to-end systems which involve clustering steps.
@@
RandFC-L-MLP 21.71 11.60 15.89 21.38
NEST-L-MLP 16.83 7.88 11.71 20.22
RandFC-L-Sortformer-HL [37] 18.93 9.39 13.56 20.15
NEST-L-Sortformer-HL [37] 16.06 6.40 11.02 16.85
@@ III-E Results on Speaker Diarization
To assess the impact of NEST on speaker diarization, we train two variants of end-to-end diarization models: (1) simple two-layer multi-layer-perceptron (MLP) on top of FastConformer encoder and train with PIL; (2) more sophisticated Sortformer [37] hybrid loss (HL) model with post processing (PP) [^4] Post-processing parameters were tuned separately for DIHARD3 and CALLHOME on corresponding training parts. with 18 layers of transformer on top of the encoder. We also apply NEST and random initialization to both models for comparison. For training data, we use a combination of 2030 hours of real data (Fisher English [24], AMI Mix-Headset train+dev [48], ICSI [49], DIHARD3 dev [50], VoxConverse v0.3 [51], AISHELL-4 [52], CALLHOME-part1555We follow splits from the Kaldi x-vector recipe [53] by using part1 for training and part2 for evaluation. [54]) and 5150 hours of simulated data (composed from LibriSpeech [43] and SRE [55, 56]) generated by the NeMo speech data simulator [57]. We evaluate models’ performance on DIHARD3-eval [50] and CALLHOME-part2 [54].